OpenStudy (anonymous):

Hannah has a jar filled with 150 coins made up of quarters and nickels. The total value of the coins is $28.30 If x represents the number of quarters in the jar, what is an equation that can be used to find the number of quarters? &How many quarters and how many nickels are there?

OpenStudy (hoblos):

sorry but can you tell me nickels=how many $

OpenStudy (anonymous):

uhm it doesnt tell me. these are the answer choices though. A) 0.05x + 0.25(150 x) = 28.30 B) 0.25x + 0.05(x 150) = 28.30 C) 0.25x + 0.05(150 x) = 28.30 D) 0.05x + 0.25(x 150) = 28.30 & A) 46 quarters; 104 nickels B) 75 quarters; 75 nickels C) 104 quarters; 46 nickels D) 95 quarters; 55 nickels

OpenStudy (hoblos):

C

OpenStudy (hoblos):

but there is a <-> between 150 x

OpenStudy (anonymous):

so C for both questions?

OpenStudy (hoblos):

yup

OpenStudy (anonymous):

ok thank you!

OpenStudy (hoblos):

no no it is A for the 2nd

OpenStudy (anonymous):

oh ok

OpenStudy (hoblos):

welcome /^_^\